OpenStudy (anonymous):

The average of four different positive integers is 9. What is the greatest value for one of the integers?

OpenStudy (anonymous):

16?

OpenStudy (uri):

x+x+x+x=9 ----------- 4 4x=36 x=36/4? not sure..

OpenStudy (anonymous):

The 4 integers have to add up to 36; Since we are trying to find the largest one of the four, we try to minimize the other 3. To minimize the other 3, we choose the smallest positive integers, which are, 1,2,3 Since all 4 have to add up to 36, the equation representing this statement is \[1+2+3+x=36\]

OpenStudy (anonymous):

So the largest that any one of the numbers could be is 30